OpenStudy (anonymous):

cos x= root 2 over 2, x= pi/4 use substitution to determine whether the given x-value is a solution of the equation.

OpenStudy (anonymous):

\[\cos (x)=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

\[\cos (\Pi/4)=\frac{ \sqrt{2}}{ 2 }\]

OpenStudy (anonymous):

Yeah, that's what I wrote, but how do I determine if it is a solution or not?

OpenStudy (anonymous):

so what is cos of pi/4?

OpenStudy (anonymous):

yes it is a solution

OpenStudy (anonymous):

Okay thanks. How were you able to know that it is a solution? What do I need to do?